Category Archives: HETDEX

The is our Dark energy experiment with VIRUS

The week of October 2nd

This week we are pleased to announce that a new VIRUS unit was installed in side two of the VIRUS enclosure. This brings us to 22 VIRUS units or 44 spectrographs. We also took a little time in the last engineering run to add on some valves to the vacuum fittings which will allow us to cold pump on the VIRUS units which takes far less time to do than to warm up and then repump which was our older methodology. Keeping 22 VIRUS units going is starting to be a little easier but still takes a lot of management.

In addition to the work on VIRUS, we have also installed in the coherent fiber bundles for the HPF. These coherent fiber bundles will be used to setup stars on HPF science fibers. HPF will arrive in the coming weeks and we are very excited to get our first high resolution instrument on sky in the coming months.

Week of April 2nd

The next few days are going to be very busy at the HET. The HET team along with help from Austin will be installing IFU’s for the second virus enclosure in preparation for more virus units to be installed. This involves stringing fiber cables from the enclosure up to the top of the tracker to be attached to the IHMP (Input Head Mounting Plate). The team will also be investigating and trying to remedy problems with the IHMP dither mechanism. During this time there will be no science or engineering related night work due to immobility of tracker, structure or dome for the IFU deployment.

Week of December 19th

This week has been an exciting week ramping up to science.  The Board of Visitor Staff Excellence Award Winners were announced by Director Armandroff and are as follows: Henry Cantu, Angela Davis, Steve Odewahn, and Trent Peterson.  Congratulations to the winners.

We are quickly moving in to science mode with the upcoming new moon.  Currently this week the first half of each night is being used primarily for science with the second half used for engineering due to the moon.  The weather looks to be clearing over the holiday weekend, so hopefully we will be able to collect some good data for the HET community!

Week of September 19th

Big Thunderstorm

One of the many powerful thunderstorms that passed by us this month. This one produced hail just north of the HET.

This week we finally had a break in the stormy weather and started wave front sensor commissioning.  A team from Austin came out and installed an imager and wave front sensor in the IHMP (Input Head Mounting Plate) to calibrate all other wave front sensors on the telescope.  These devices once calibrated will help keep the tracker in perfect alignment with the mirror to optimize the observations we are taking for science.  This will be the main focus for the next month.  Hopefully the weather is more cooperative than it has been.  If you would like to see a time-lapse of some of the recent storms follow this link: https://www.youtube.com/watch?v=shRD2WmPWd4

The week of August 15th

Virus and Controllers

16 Virus units and their controllers lined up in the HET receiving bay. The controllers are the small boxes near the center of the image.

This week has been a busy one for the virus units at the HET.  All of the virus units were removed from the enclosure, backfilled, pumped to vacuum and reinstalled.  While doing this cooling fans were installed on all of the virus controllers to help keep the electronics cool.  They hope that this will make the virus system a bit more stable.

The HET has been churning out lots of science during dark time this month as well.  All groups that have submitted targets have received data!

The week of Nov 4

In the past week we finished cleaning up the upper hex. In the process we had to move some of the cameras around in our internal views of the dome so the view located at:

http://het.as.utexas.edu/wfu might look a bit strange until we get around to moving the cameras back into place.

This week we will be starting the metrology of the upper hex in preparation for putting on the upper and lower X rails. This metrology is done with a laser tracker and small “nest” mirrors placed around the telescope structure. Other activities this week will include finishing up the cable-ways in the lower electrical room and spectrograph room, work on the structure azimuth cable wrap, and work on the static cable ways. It is amazing the amount of power wires, optical cables, glycol and air lines that need to be strung up.

We are still right on schedule.

The week of Oct 28

In the past week we finished most of the work on the Virus Support Structure (VSS) and on Friday were able to have the first rotation including the new air bearings under the VSS. In addition, we had some of the team working on sealing the dome enclosure. Vicki worked on the outer vents at the base of the ring wall and Logan and Kathryn worked on the dome and sealed the nodes near top.

DomeNodes1

This week we will be finishing the clean up of the upper hex removing the trollies, encoders and last of the interconnects. We will also make room for the new JLG man-lift (we call Jr.) on the dome floor. This work is in preparation for the work that John Good will start in Nov.

The week of Oct 21

This past week we worked on getting the Virus Support Structure (VSS) cross supports and over-turn preventers welded into place. Other activities included test digs around the location of the liquid nitrogen tank and the vault for the remote thermal area, setting up the air system for the VSS and continued work on the cable-ways for the lower electrical room.

This next week we should finish the VSS work which means that we should be able to unlock structure rotation. If all goes well we will now have the Structure firmly welded to the VSS and they should rotate together as if they were one system.

The week of Oct 14

In the past week we had great guys from Crane Services out to the HET to remove the old tracker bridge from the dome, move the new lower X and upper X assemblies onto the dome floor, and move the VSS into place. It all went very smoothly and they finished on Thursday about 1 hour before a small windstorm (by West Texas standards) moved in.

There are numerous pictures taken by:

Jerry Martin

Dave Doss of the Old Tracker lift

Dave Doss of the VSS installation

and here is a picture of the tracker removal taken from Renny Spencer who was purposefully stranded up on the upper X beam to monitor the lift.

310

There is even a movie produced by Eusebio Terrazas:

http://het.as.utexas.edu/wfu/

This next week we will be tuning up the VSS, putting in the diagonal bracing and getting ready to hook it up to the air bearing system. Other work at the HET includes putting in cableways in the lower electrical room and spectrograph room, stripping the old coating from the primary mirror segments that were removed before the upgrade began, and applying extra sealant to the dome nodes.

The week of Oct 7

Last week we finished the painting of the CCAS tower and the removal of the equipment on upper X hex.

This week we will expect a crane to arrive on Monday and from Tuesday through Friday we will be removing the old tracker bridge, install the Virus Support Structure and stage the upper and lower X rail assemblies for the new tracker onto the dome floor. The upper and lower X rail assemblies will be shipped from CEM in Austin to McDonald on Monday so that they will be on site Tuesday morning. The road to the top of the mountain will be closed to non-essential personnel.

We are still on schedule.